OpenStudy (anonymous):

can anyone walk me through finding out X when -3+2x=-x+6?

OpenStudy (anonymous):

- 3 +2x = -x + 6 -3 -6 + 2x = -x + 6 - 6 -9 + 2x = -x -9 +2x - 2x = -x -2x -9 = -3x divide both sides by -3 3 = x